OpenStudy (anonymous):

HELP MEEEE !! Find the slope of the secant to the curve f(x)=0.75(3)^x -1 between: x=o and x=1

OpenStudy (kc_kennylau):

When x=0, what's the value of f(x)? When x=1, what's the value of f(x)?

OpenStudy (anonymous):

what do you do after you plug in the numbers?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

You get two points, (0, ??) and (1, ??).

OpenStudy (kc_kennylau):

Then you can calculate the slope using the formula \(\huge m=\dfrac{y_2-y_1}{x_2-x_1}\)

OpenStudy (anonymous):

okay so for part b is it the same thing?

OpenStudy (anonymous):

Extend the results from part a) to determine the slope of the tangent to the curve at x = 0, accurate to 3 decimal places.

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

sorry but which grade are you in? :)

OpenStudy (anonymous):

12

OpenStudy (kc_kennylau):

Do you know calculus?

OpenStudy (anonymous):

I dont have it yet

OpenStudy (kc_kennylau):

I think this is what the teacher wants you to do (Just I think): So the tangent is just the secant between x=0 and x=0.000000000000000001

OpenStudy (anonymous):

oh ok

OpenStudy (kc_kennylau):

ofc you don't need so many 0s

OpenStudy (anonymous):

thanks :)

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

An antique vase was appraised in 2000 for $8000. If the vase appreciates in value by 6% per year, what is its estimated value in the year 2040, to the nearest thousand dollars?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

\[\huge F=I(1+r\%)^t\] A=Final value P=Initial value r=rate y=time

OpenStudy (anonymous):

what would be the answer then?

OpenStudy (kc_kennylau):

\[8000\times(1+6\%)^{40}\]

OpenStudy (anonymous):

so the answer is 82285??

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

The nearest THOUSAND dollars

OpenStudy (anonymous):

83000

OpenStudy (kc_kennylau):

And if anything it's 82286 not 82285

OpenStudy (kc_kennylau):

nope

OpenStudy (anonymous):

82000???

OpenStudy (kc_kennylau):

exactly

OpenStudy (anonymous):

oh thanks!!

OpenStudy (anonymous):

The world population doubles approximately every 40 years. If the population was approximately 4.5 billion in 1980, what is the projected approximate population in the year 2080? @kc_kennylau

OpenStudy (kc_kennylau):

How many years are there between 1980 and 2080?

OpenStudy (anonymous):

100

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

How many years is a period?

OpenStudy (anonymous):

12??

OpenStudy (anonymous):

idk

OpenStudy (kc_kennylau):

40

OpenStudy (anonymous):

ohhh

OpenStudy (kc_kennylau):

How many periods are there?

OpenStudy (anonymous):

40

OpenStudy (anonymous):

100

OpenStudy (kc_kennylau):

40 years=1 period, 100 years=? periods

OpenStudy (anonymous):

2 and something

OpenStudy (kc_kennylau):

2.5

OpenStudy (anonymous):

so can you put it into the equation for me ?

OpenStudy (kc_kennylau):

Now plug all the values into the formula: \[\huge F=I(1+r\%)^p\] F=Final value (The number of ppl in 2080) I=Initial value (The number of ppl in 1980) r=Rate (By how many percent does it increase in one period?) p=number of Periods (How many periods are there?)

OpenStudy (anonymous):

what is r?

OpenStudy (kc_kennylau):

100

OpenStudy (kc_kennylau):

Since doubling means increasing by 100%

OpenStudy (anonymous):

answer is 461334.5??

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

nope, how did you get this value?

OpenStudy (anonymous):

by plugging in the numbers

OpenStudy (anonymous):

4.5(1+100%)^2.5 ???

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

yep

OpenStudy (anonymous):

thats what I got :( what did you get ???

OpenStudy (kc_kennylau):

I got 25 billion...

OpenStudy (anonymous):

ohhh I see what I did wrong

OpenStudy (kc_kennylau):

so you get it now? :)

OpenStudy (anonymous):

Inflation is causing a rise of approximately 2% per year in the cost of living. A bag of milk costs $3.95 now. Estimate the cost five years from now. A movie ticket costs approximately $11.00 now. If inflation continues at 2% per year, when will the ticket cost $13.50? If the move ticket costs approximately $11.00 now, how long ago did it cost $4.00?

OpenStudy (anonymous):

Inflation is causing a rise of approximately 2% per year in the cost of living. (6 marks) a.A bag of milk costs $3.95 now. Estimate the cost five years from now. b. A movie ticket costs approximately $11.00 now. If inflation continues at 2% per year, when will the ticket cost $13.50? c. if the move ticket costs approximately $11.00 now, how long ago did it cost $4.00?

OpenStudy (kc_kennylau):

Can you try to do it yourself first? :)

OpenStudy (anonymous):

same equation right?

OpenStudy (kc_kennylau):

of course :)

OpenStudy (anonymous):

but Im confused :( I suck at this

OpenStudy (anonymous):

can you put it in the equation and then I can solve it ...

OpenStudy (kc_kennylau):

For part a: 1. What are you trying to find, i.e. the unknown? 2. What values are given? 3. Can you put it in the equation? 4. Can you solve the equation? For part b: 1. What are you trying to find, i.e. the unknown? 2. What values are given? 3. Can you put it in the equation? 4. Can you solve the equation? For part C: 1. What are you trying to find, i.e. the unknown? 2. What values are given? 3. Can you put it in the equation? 4. Can you solve the equation? Sorry, I cannot just do it for you, since I discover that this method cannot help you learn. Just doing the homework for others may be able to help some people, but not all people.

OpenStudy (anonymous):

@kc_kennylau for a I got 22

OpenStudy (anonymous):

is t right ??

OpenStudy (kc_kennylau):

nope :/

OpenStudy (anonymous):

I plugged it into the equation :(

OpenStudy (kc_kennylau):

can you show me how you plugged? :)

OpenStudy (anonymous):

3.95(1+100)^2.5

OpenStudy (kc_kennylau):

the rate is not 100% this time, and the number of periods ain't 2.5 this time...

OpenStudy (anonymous):

it will be 1

OpenStudy (kc_kennylau):

"a rise of approximately 2% per year" "five years from now"

OpenStudy (kc_kennylau):

Sorry give me two minutes, I have to reset my modem :/

OpenStudy (kc_kennylau):

hi i'm back :)

OpenStudy (anonymous):

hii

OpenStudy (anonymous):

so what is r?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

2

OpenStudy (anonymous):

and the period?

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

5

OpenStudy (anonymous):

4.36

OpenStudy (kc_kennylau):

yep

OpenStudy (kc_kennylau):

you still can't get the answer by yourself... what do you not understand? :)

OpenStudy (anonymous):

idk just the period

OpenStudy (anonymous):

for part b we're solving p ?

OpenStudy (kc_kennylau):

yep

OpenStudy (kc_kennylau):

so the period is the thing after "per", like if it says "per year", the period is 1 year. and the number of period is the number of years divided by the length of one period. For example, if the question asks for 5 years, the number of period is 5.

OpenStudy (anonymous):

so what is the p for part b?